Suppose your firm is considering investing in a project with cash flows below, that the required rate of return on projects of this risk class is 8%, and that the maximum allowable payback and discounted payback statistics for the project are 3.5 and 4.5 years respectively.

Time 0. 1. 2. 3. 4. 5. 6.

Cashflow -$5,000 $1,200 $2,400 $1,600 $1,600 $1,400 $1,200

Use the payback decision rule to evaluate this project; should it be accepted or rejected?

Please Use Excel while showing work
Answer #1

Payback period is the time taken for the cumulative cash flows to equal zero

Payback period = 2 + (cash flow required in time 3 for cumulative cash flows to equal zero / time 3 cash flow) = 2 + (1400 / 1600) = 2.875 years

The maximum allowable payback period for the project is 3.5 years.

The project should be accepted as the project's payback period is lower than the maximum allowable payback period.

1 A B C Cumulative Time Cash Flow Cash Flow 0 $ (5,000) $ (5,000) 1 $ 1,200 $ (3,800) 2 $ 2,400 $ (1,400) 3 $ 1,600 $ 200 4 $